OpenStudy (anonymous):

Dillon has a bag of marbles. There are 4 blue, 1 green, 1 white, 3 black, and some red marbles in the bag. If the probability of pulling out a red marble is 3/4, how many red marbles are in the bag?

OpenStudy (anonymous):

\[\frac{x}{9+x}=\frac{3}{4}\]

OpenStudy (anonymous):

i think its 27 total red marbles??

OpenStudy (anonymous):

27 marbles